sue bought 4 pounds of coffee and two pounds of tea for $31.50 last week. this week she bought 3 pounds of coffee and 4 pounds of tea for $38.25. what is the cost per pound of both the coffee and tea?
Answer by checkley77(12844)

4c+2t=31.50 multiply this equation by -2 & add.
3c+4t=38.50
-8c-4t=-63
----------------------
-5c=-24.5
c=-24.5/-5
c=$4.90 price for a pound of coffee.9
4*4.90+2t=31.50
19.6+2t=31.50
2t=31.50-19.6
2t=11.90
t=11.90/2
t=$5.95 price for a pound of tea.
Proof:
3*4.90+4*5.95=38.50
14.70+23.8=38.50
38.5=38.5